Circular No. 8883

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V458 VULPECULAE D. K. Lynch, R. W. Russell, and R. J. Rudy, The Aerospace Corpora-

tion; and C. E. Woodward, University of Minnesota, report 0.8- to 5.5-μm spectroscopy of V458 Vul (cf. IAUC 8863, 8878) on Oct. 10 and 11 UT using the Infrared Telescope Facility (+ SpeX): “V458 Vul has the strongest lines of neutral helium that we have encountered in a nova. He I 1083.0-nm has an equivalent width of > 1400 nm and is 30 times larger than Paβ. Similarly, He I 2058.1-nm is seven times as strong as Brγ. Numerous other lines of He I are present. The lines of both helium and hydrogen are flattopped and have FWHM of 2700 km/s. The O I lines are still present and indicate a reddening of E(B − V) = 0.6. He II features are just beginning to emerge, so the He I features will likely weaken slightly in the future as more helium becomes doubly ionized. No thermal emission from dust was detected.”

COMETS C/2007 K19, C/2007 K20, C/2007 L1–L6 (SOHO) Further to IAUC 8882, additional Kreutz sungrazing comets have been

found on SOHO website images. C/2007 K19 and L5 were extremely faint (mag ∼ 8.5) and diffuse, the former being also elongated. C/2007 K20 and L6 were very faint (mag ∼ 8) and diffuse. C/2007 L1 was stellar in appearance in C3 images, and slightly diffuse (mag ∼ 7.5) with no tail in C2 images. C/2007 L2 was stellar in appearance in C3 images, and was of mag ∼ 5.5 with an extremely faint, thin tail in C2 images. C/2007 L3 peaked at mag ∼ 3 with a dense, wide tail ∼ 45′′ long on June 8.171 UT; while the comet’s “head” disappeared at ∼ June 8.181, the tail was still visible at June 8.917 (when it had extended into a diffuse arc of apparent length around 1^{o}.5), and this comet was also visible in coronagraphs on both STEREO spacecraft. C/2007 L4 was stellar in appearance (mag ∼ 8).

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 K19 May 28.326 4^{h}17.^{m}0 +19^{o}32′ C2  RK 2007-T93
C/2007 K20 31.129 4 29.8 +20 02 C2 TH 2007-T93
C/2007 L1  June 4.305 4 46.4 +20 13 C3/2 BZ 2007-T94
C/2007 L2  4.971 4 52.1 +18 39 C3/2 KB 2007-T94
C/2007 L3  6.388 4 56.3 +18 10 C3/2 SY 2007-T94
C/2007 L4  7.704 4 59.0 +20 53 C2  BZ 2007-T94
C/2007 L5  9.371 5 03.7 +20 56 C2  AW 2007-T94
C/2007 L6  10.621 5 08.4 +21 06 C2  BZ 2007-T117
2007 October 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green